### v2

# Harnessing The Collective Wisdom: Fusion Learning Using Decision Sequences From Diverse Sources

Abstract
Learning from the collective wisdom of crowds is related to the statistical notion of fusion learning from multiple data sources or studies. However, fusing inferences from diverse sources is challenging since cross-source heterogeneity and potential data-sharing complicate statistical inference. Moreover, studies may rely on disparate designs, employ myriad modeling techniques, and prevailing data privacy norms may forbid sharing even summary statistics across the studies for an overall analysis. We propose an Integrative Ranking and Thresholding (IRT) framework for fusion learning in multiple testing. IRT operates under the setting where from each study a triplet is available: the vector of binary accept-reject decisions on the tested hypotheses, its False Discovery Rate (FDR) level and the hypotheses tested by it. Under this setting, IRT constructs an aggregated and nonparametric measure of evidence against each null hypotheses, which facilitates ranking the hypotheses in the order of their likelihood of being rejected. We show that IRT guarantees an overall FDR control if the studies control their respective FDR at the desired levels. IRT is extremely flexible, and a comprehensive numerical study demonstrates its practical relevance for pooling inferences. A real data illustration and extensions to alternative forms of Type I error control are discussed.
Keywords: Crowdsourcing, E-values, False Discovery Rate, Integrative inference, Meta-analysis
1 Introduction
Learning from the wisdom of crowds is the process of synthesizing the collective wisdom of disparate participants for performing related tasks, and has evolved into an indispensable tool for modern scientific inquiry. This process is also known as ‘Crowdsourcing’, which has generated tremendous societal benefits through novel drug discovery (Chodera et al., 2020), new machine learning algorithms (Sun et al., 2022), product design (Jiao et al., 2021) and crime control (Logan, 2020), to name a few. Harnessing such collective wisdom allows incorporating diverse opinions as well as expertise into the decision-making process and ultimately broadens the transparency of scientific findings (Surowiecki, 2005).
Synthesizing the collective wisdom of crowds is related to the statistical notion of fusion learning from multiple data sources or studies (Liu et al., 2022; Guo et al., 2023). Over the past decade, techniques for such fusion learning have found widespread use for incorporating heterogeneity into the underlying analysis and increasing the power of statistical inference. For instance, in recent years the accrescent volume of gene expression data available in public data repositories, such as the Gene Expression Omnibus (GEO) (Barrett et al., 2012) and Array Express (Rustici et al., 2012), have facilitated the synthesis and reuse of such data for meta-analysis across multiple studies. Biologists can use OMiCC (Shah et al., 2016), a crowdsourcing web platform, for generating and testing hypotheses by integrating data from diverse studies. Relatedly, sPLINK (Nasirigerdeh et al., 2022), a hybrid federated tool, allows conducting privacy-aware genome-wide association studies (GWAS) on distributed datasets. In distributed multisensor systems, such as wireless sensor networks, each node in the network evaluates its designated region in space by conducting simultaneous test of several hypotheses. The decision sequence is then transmitted to a fusion center for processing. Sensor fusion techniques are used to synthesize the data from multiple sensors, which provides improved accuracy and statistically powerful inference than that achieved by the use of a single sensor alone (Hall and Llinas, 1997; Shen and Wang, 2001; Jamoos and Abuawwad, 2020).
However, fusing inferences from diverse sources [^1] We will use the terms ‘data-source’, ‘study’ and ‘nodes’ interchangeably throughout this article. is challenging for several reasons. First, cross-source heterogeneity and potential data-sharing complicate statistical inference. For instance, in microarray meta-analysis, an integrative false discovery rate (FDR) analysis of the multiple hypotheses often requires making additional assumptions such as study independence and strong modeling assumptions to capture between-study heterogeneity. Second, prevailing data privacy norms may forbid sharing even summary statistics across the studies for an overall analysis. This is particularly relevant in the context of genomic data where NIH (National Institutes of Health) restricts the availability of dbGaP (database of Genotypes and Phenotypes) data to approved users (Couzin, 2008; Zerhouni and Nabel, 2008). Also, in wireless sensor networks, communication limits may restrict sharing all information from the nodes to the fusion center. Third, in the case of meta-analysis where fusion learning is widespread, different studies may rely on disparate designs and widely different modeling techniques for individual inferences. Besides introducing algorithmic randomness into the individual inferences, it is also unclear how such inferences can be integrated and subsequently interpreted. Fourth, often significant statistical and computational experience is required for integrating the individual inferences. This may preclude investigators without substantial statistical training to perform such analyses.
In this paper, we develop a framework for fusion learning in multiple testing that seeks to overcome the aforementioned challenges of integrative inference across multiple data sources. Our framework, which we call IRT for Integrative Ranking and Thresholding, operates under the setting where from each study a triplet is available: the study-specific vector of binary accept / reject decisions on the tested hypotheses, the FDR level of the study and the hypotheses tested by the study. Under this setting, the IRT framework consists of two key steps: in step (1) IRT utilizes the binary decisions from each study to construct nonparametric evidence indices which serve as measures of evidence against the corresponding null hypotheses, and in step (2) the evidence indices across the studies are fused into a discriminatory measure that ranks the hypotheses in the order of their likelihood of being rejected. The proposed fusion learning framework has several distinct advantages. First, the IRT framework guarantees an overall FDR control as long as the individual studies control the FDR at their desired levels. This FDR control holds under arbitrary dependence between the fused evidence indices from step (2). Second, IRT is extremely simple to implement and is broadly applicable without any model assumptions. This particular aspect is especially appealing because IRT synthesizes inferences from diverse studies irrespective of the underlying multiple testing algorithms employed by the studies. Third, the evidence indices in our framework are closely related to “$e-$values” (see Shafer (2021); Vovk and Wang (2021); Grünwald et al. (2023); Ramdas et al. (2020); Wasserman et al. (2020) for an incomplete list) for hypothesis testing. Besides being a natural counterpart to the popular $p-$values in statistical inference, $e-$values are relatively more robust to model misspecification and particularly to dependence between the $p-$values. In our numerical experiments we find that when the $p-$values are dependent, IRT is substantially more powerful than methods that rely on a conversion from $p-$values to $e-$values for pooling inferences. Additionally, when the distribution of the test statistic under the null hypothesis is misspecified and the corresponding $p-$values are sub-optimal, procedures that rely directly on the magnitude of such $p-$values for pooling inferences may fail to control the FDR. In contrast, IRT is relatively robust to such misspecification in our simulation studies. For almost a century, $e-$values have been used in Statistics, often disguised as likelihood ratios, Bayes factors, stopped nonnegative supermartingales. See Ramdas et al. (2022) for a survey on additional settings, such as betting scores, game-theoretic statistics and safe anytime-valid inference, where $e-$values arise naturally. Finally, to the best of our knowledge, IRT is the first fusion learning framework for multiple testing that relies on the binary decision sequences which are relatively more private than summary statistics. We rely on a novel construction of $e-$values from the accept-reject decisions which are also related to the all-or-nothing bet of Shafer (2021) for testing a single null hypothesis.
Figure 1: A pictorial representation of the IRT framework. In panel (1) $d$ studies are testing a study-specific subset of $m$ null hypotheses. Here the black solid squares represent hypotheses that are not tested by the individual studies. In panel (2), these studies use different multiple testing algorithms that are designed to control their respective FDR at level $\alpha_{j},~{}j=1,\ldots,d$. The ‘X’-marked squares in panel (3) depict the rejected hypotheses and represent the data that is available to the IRT framework. It includes the binary decision vectors $\delta_{ij}$ from each study, the study-sepcific FDR level $\alpha_{j}$ and the subset of $m$ hypotheses evaluated by each study. In panel (4), the IRT framework converts each binary decision into a nonparametric evidence index such that a higher index, displayed in a darker shade of red, represents a larger likelihood for rejecting that study-specific null hypothesis. In panel (5), the study-specific evidence indices are fused in a manner such that the aggregated indices rank the $m$ hypotheses and, as demonstrated in panel (6), the rankings can be used to determine a cutoff for valid FDR control at a pre-determined level $\alpha$.
1.1 Outline of the proposed approach
Figure 1 provides a pictorial representation of the IRT framework under a simplified setting. A precise formulation and formal mathematical statements are deferred until Section 2. Panel (1) of Figure 1 represents a scenario where $d$ studies are testing a study-specific subset of $m$ null hypotheses. Here the black solid squares represent hypotheses that are not tested by the individual studies. In Panel (2), these individual studies use different multiple testing algorithms that are designed to control their respective FDR at level $\alpha_{j},~{}j=1,\ldots,d$. The ‘X’-marked squares in Panel (3) depict the rejected hypotheses and represent the data that is available to the IRT framework for synthesis. It includes the binary decisions $\delta_{ij}$ from each study, the study-specific FDR level $\alpha_{j}$ and the subset of $m$ hypotheses evaluated by each study. Panels (4)-(6) illustrate the IRT framework. Particularly, in Panel (4), the IRT framework converts each binary decision $\delta_{ij}$ into a nonparametric evidence index such that a higher index, displayed in a darker shade of red, represents a stronger conviction against that study-specific null hypothesis. In Panel (5), the study-specific evidence indices are fused in a manner such that the aggregated indices rank the $m$ hypotheses and, as demonstrated in Panel (6), the rankings can be used to determine a cutoff for valid FDR control at a pre-determined level $\alpha$. While Figure 1 introduces IRT under a setting where $d$ studies are testing a study-specific subset of $m$ null hypotheses, it is much more broadly applicable, including the setting where pooling inferences for a common set of $m$ hypotheses across the $d$ studies is of interest. Section 2 presents those details.
1.2 Related literature and main methodological contributions
The IRT framework is related to multiple strands of literature, which includes FDR control using $e-$values, integrative multiple testing under data-privacy and algorithmic derandomization. We discuss how IRT differs from these existing body of works.
Multiple testing with $e-$values - Recently, there has been a proliferation of interest in developing methods for simultaneous hypotheses testing using $e$-values. See, for instance, Wang and Ramdas (2022); Ignatiadis et al. (2024); Chi et al. (2022); Vovk and Wang (2023); Xu and Ramdas (2023) and the references therein. In these works, the $e-$values are typically constructed from either $p-$values or likelihood ratios. In contrast, the generalized $e-$values (Wang and Ramdas, 2022; Ren and Barber, 2024) in our setting arise from the accept / reject decisions of the corresponding null hypotheses and are constructed in a nonparametric fashion.
Integrative multiple testing under data-privacy - The literature on the evolving field of integrative multiple testing under data-privacy is also related to our work. For instance, under the restricted data-sharing mechanism of Wolfson et al. (2010), Liu et al. (2021) propose an integrative high-dimensional multiple testing framework with asymptotic FDR control. Our work differs from Liu et al. (2021) in two main aspects. First, we consider a relatively more stringent data-privacy regime where the studies are only allowed to share their decision sequences, the FDR level and the hypotheses tested by them. Second, the integrative FDR control offered by IRT is non-asymptotic if the individual studies guarantee a non-asymptotic control of their respective FDR levels (see Section 4.2). This aspect is particularly appealing when the underlying multiple testing problem is small-scale.
Algorithmic stability - Numerous methods have been proposed for mitigating algorithmic randomness in recent years. For instance, Bashari et al. (2023) propose a method to aggregate conformal tests for outliers obtained with repeated splits of the same data set, while controlling the FDR. In the context of high dimensional variable selection, Ren and Barber (2024) develop a methodology for derandomizing model-X knockoffs (Barber and Candès, 2015; Candes et al., 2018) with provable FDR control while Dai et al. (2023a, b) develop a multiple data-splitting method with asymptotic FDR control to stabilize variable selection results. The IRT framework can be similarly viewed as a method for mitigating algorithmic randomness that may potentially arise from the use of myriad multiples testing procedures for FDR control by the $d$ studies. Crucially, however, our framework does not rely on hard-to-verify assumptions for FDR control, such as Dai et al. (2023a, b), and in contrast to Ren and Barber (2024); Bashari et al. (2023); Li and Zhang (2023), IRT only requires access to the binary decision vector of each study for integrative FDR control. In fact, in Section 4.1 we show that several recent works that leverage $e-$values for aggregating and derandomizing specific FDR methods (see Li and Zhang, 2023; Ren and Barber, 2024; Zhao and Sun, 2024) are special cases of our evidence index (Equation (1)) from an asymptotic perspective.
The ensuing discussion summarizes our main methodological contributions.
  1. 1.
    Construction of evidence indices from binary decisions - at an intuitive level, a multiple testing rule essentially translates evidence indices, such as $p-$values, $e-$values, $z-$scores, into binary decisions. This paper presents a comprehensive solution to the “reverse testing problem”, which involves converting binary decisions into evidence indices. We choose $e-$value as an evidence index because of its many desirable theoretical properties as pointed out in Wang and Ramdas (2022); Vovk and Wang (2021). While the conversion of binary decisions to $e-$values has been explored in Ren and Barber (2024); Li and Zhang (2023), these studies are limited to specific testing processes, such as knockoffs and BH, and focus solely on testing rules that guarantee FDR control. In contrast, we show how $e-$values can be constructed from any testing procedure that guarantees some kind of type I error control, such as FDR, $k-$FWER, and PFER (see Section 7).
  2. 2.
    Evidence aggregation via multiplication - for each hypothesis, IRT fuses the evidence indices via their arithmetic mean. However, when inferences are pooled with the goal of higher reliability then such an aggregation scheme often renders IRT extremely conservative. In Section 3.2, we provide precise conditions under which the evidence indices can be fused using a product rule. These conditions are related to the notion of partial exchangeablilty on the set of nulls, commonly used in the conformal inference literature (Bates et al., 2023; Liang et al., 2024), and symmetric decision rules, which arise naturally in conventional settings where all hypotheses undergo simultaneous testing without the inclusion of auxiliary side information. The underlying inference pooling framework that relies on this product aggregation scheme is called IRT and is, typically, more powerful than IRT.
  3. 3.
    A flexible framework for pooling inferences - while the proposed framework is extremely simple, it enjoys substantial flexibility with which it can be used in myriad applications. For instance, IRT(IRT) is applicable when only some of the studies share $p-$values, while for the others only the binary decisions are available. In such a setting, IRT (IRT) can be used in conjunction with a p2e calibrator, such as those proposed in Vovk and Wang (2021), for pooling inferences (see Remark 3). Furthermore, IRT (IRT) can pool inferences across studies for an overall FDR control even when studies employ different type I error control metrics, such as family-wise error rate or per-family error rate control (see Remark 4). Finally, in many practical settings prior knowledge may dictate that some studies share data amongst themselves or rely on auxiliary side information for multiple testing. In this scenario, a hybrid scheme IRT H, that combines IRT and IRT, can synthesize inferences across all the studies and ensure valid FDR control (see Remark 1).
  4. 4.
    Robustness to misspecified null distribution - when the summary statistics, such as $p-$values, are available from all studies then neither IRT nor IRT are natural candidates for pooling inferences as they may lead to power loss. This is indeed true as long as the null distribution of the test statistic is correctly specified. The numerical experiments in Section 6 reveal that when the null distribution is misspecified, methods that rely directly on the magnitude of the $p-$values, may fail to control FDR at the desired level. In contrast IRT, IRT and IRT H are relatively robust to such misspecification. This is an important practical setting since typically the null distribution specification relies on large sample approximations, and such large samples may not be available for some studies.
1.3 Organization
The article is laid out as follows: Section 2 presents the IRT framework for integrative FDR control. In Section 3 we present IRT that relies on an alternative scheme for evidence aggregation. Additional insights about the IRT framework are discussed in Section 4. A real data illustration is presented in Section 5 while Section 6 reports the empirical performance of IRT on synthetic data. Extensions to alternative forms of Type I error control are discussed in Section 7. The article concludes with a summary in Section 8. The data and R-code used in this article are available at https://github.com/trambakbanerjee/IRT.
Notations
Let =$\mathcal{M}=\{1,\dots,m\}$ and consider testing $m$ null hypotheses $H_{01},\dots,H_{0m}$. Denote $\mathcal{H}_{0}=\{i:H_{0i}\ \text{is true}\}$ and $\mathcal{H}_{1}=\mathcal{M}\setminus\mathcal{H}_{0}$, respectively, as the set of true null and non-null hypotheses. Here $\setminus$” denotes the usual set difference operator. Let $\mathbb{I}(\cdot)$ be the indicator function that returns 1 if the condition is met and 0 otherwise, and denote $\|\mathbf{w}\|_{p}$ as the $\ell_{p}$-norm of vector $\mathbf{w}$. In the sequel, $a\vee b$ will denote $\max\{a,b\}$ for two real numbers $a$ and $b$, $[d]=\{1,\ldots,d\}$ and the cardinality of a set of positive integers $\mathcal{I}$ will be denoted by $|\mathcal{I}|$.
2 IRT: integrative FDR control using binary decision sequences
We consider a setting where the $\mathcal{M}$ hypotheses are tested by $d$ studies with study $j$ testing $\mathcal{M}_{j}$ hypotheses. Here $\mathcal{M}_{j}\cap\mathcal{M}\neq\emptyset$ and $|\mathcal{M}_{j}|=m_{j}$ for all $j\in[d]$. Let $\theta_{i}=\mathbb{I}(H_{0i}\text{~{}is false})$ be an indicator function that gives the true state of the $i$-th testing problem and denote $\delta_{ij}\in\{0,1\}$ as the decision that study $j$ makes about hypothesis test $i\in\mathcal{M}_{j}$, with $\delta_{ij}=1$ being a decision to reject $H_{0i}$. Denote the vector of decisions as $\bm{\delta}_{j}=(\delta_{1j},\cdots,\delta_{m_{j}j})\in\{0,1\}^{m_{j}}$. A selection error, or false positive, occurs if study $j$ asserts that $H_{0i}$ is false when it actually is not. In multiple testing problems, such false positive decisions are inevitable if we wish to discover interesting effects with a reasonable power. Instead of aiming to avoid any false positives, a practical goal is to keep the false discovery rate (FDR, Benjamini and Hochberg, 1995) small, which is the expected proportion of false positives among all selections,
$\text{FDR}(\bm{\delta}_{j})=\mathbb{E}\left[\text{FDP}(\bm{\delta}_{j})\right]~{}\text{where FDP}(\bm{\delta}_{j})=\dfrac{\sum_{i\in\mathcal{M}_{j}}(1-\theta_{i})\delta_{ij}}{\max\Bigr{\{}\sum_{i\in\mathcal{M}_{j}}\delta_{ij},1\Bigr{\}}}.$
The power of a testing procedure is measured by the expected number of true positives (ETP) where,
$\mbox{ETP}(\bm{\delta}_{j})=\mathbb{E}\left[\sum_{i\in\mathcal{M}_{j}}\theta_{i}\delta_{ij}\right]$
Hence, the multiple testing problem for study $j$ can be formulated as
$\text{maximize}_{\bm{\delta}_{j}}~{}\text{ETP}(\bm{\delta}_{j})~{}\text{subject to }\mbox{FDR}(\bm{\delta}_{j})\leq\alpha_{j},$
where $\alpha_{j}\in(0,1)$ is a pre-specified cap on the maximum acceptable FDR for the $j$-th study. Our goal is to conduct inference for $\bm{\theta}=(\theta_{1},\ldots,\theta_{m})$ by fusing the evidence available in the binary decision sequences $\bm{\delta}_{j}$ provided by the $d$ studies. We develop an integrative ranking and thresholding procedure that operates on the triplet $\mathcal{D}_{j}=\{\bm{\delta}_{j},\alpha_{j},\mathcal{M}_{j}\}$ available from each study and provably controls the FDR within level $\alpha\in(0,1)$.
In the following sections we present IRT, an Integrative Ranking and Thresholding framework that involves three steps. In Step 1, IRT utilizes the binary decision sequence $\bm{\delta}_{j}$ from study $j$ to construct a measure of evidence against the null hypotheses. In Step 2, this evidence is aggregated into a discriminatory measure such that for each null hypothesis $H_{0i}$, a large aggregated evidence implies stronger evidence against $H_{0i}$. In the final Step 3, the rankings of aggregated evidence are used to determine a cutoff for FDR control. In what follows, we describe each of these steps in detail while Algorithm 1 summarizes the discussion below.
2.1 Step 1: Evidence Construction
IRT uses the information in $\mathcal{D}_{j}$ to construct an evidence index $e_{ij}$ as follows:
$e_{ij}=w_{j}\dfrac{\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}{\text{~{}where~{}}w_{j}=\frac{m_{j}}{\alpha_{j}}},\quad\forall i\in\mathcal{M}_{j}.$ (1)
The evidence weights $w_{j}$ in Equation (1) capture the relative importance of a rejection across the $d$ studies and, hence, play a key role in differentiating among them. For study $j$, the evidence weight is distributed evenly across the rejected hypotheses. Moreover, this weight is higher if study $j$ has tested more hypotheses (larger $m_{j}$) and more conservative (smaller $\alpha_{j}$). This is in perfect accordance with our intuition since the evidence needed to reject a hypothesis increases as the number of hypotheses tested increases or the target FDR level decreases. For example, if only one hypothesis is tested then rejecting that null hypothesis when the $p-$value is below $\alpha$ guarantees FDR is controlled at level $\alpha$. However, if more than one hypothesis is tested then simply rejecting null hypotheses with $p-$value at most $\alpha$ does not guarantee FDR control at level $\alpha$.
The evidence indices $\mathbf{e}_{j}=\{e_{ij}\}_{i\in\mathcal{M}_{j}}$ in Equation (1) are related to $e-$values for hypothesis testing, which serve as a natural counterpart to the widely adopted $p-$values. A non-negative random variable $e$ is an “$e-$value” [^2] We will use the notation ‘$e$’ to denote both the random variable and its realized value. if $\mathbb{E}[e]\leq 1$ under the null hypothesis. A large $e-$value provides evidence against the null hypothesis. See Vovk and Wang (2021) for a background on $e-$values for hypothesis testing. Recently, Wang and Ramdas (2022); Ren and Barber (2024) introduce the concept of generalized $e-$values which are defined as follows:
Definition 1 (Generalized $e-$values).
Let $\mathbf{e}=\{e_{1},\dots,e_{m}\}$ be a collection of random variables associated with the null hypotheses $H_{01},\dots,H_{0m}$. We say $\mathbf{e}$ is a set of generalized $e-$values if $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq m$.
Theorem 1 establishes that the evidence indices in Equation (1) are generalized $e-$values.
Theorem 1.
Suppose study $j$ controls FDR at level $\alpha_{j}$. Then, $\mathbf{e}_{j}$ is a set of generalized $e-$values associated with $\mathcal{H}_{0j}=\{H_{0i}:i\in\mathcal{M}_{j}\}$.
Proof of Theorem 1..
Based on the evidence construction in (1), we have
$\sum_{i\in\mathcal{H}_{0j}}\mathbb{E}[e_{ij}]$ $=\dfrac{m_{j}}{\alpha_{j}}\cdot\mathbb{E}\left[\dfrac{\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}\right]=\dfrac{m_{j}}{\alpha_{j}}\cdot\text{FDR}(\bm{\delta}_{j})\leq m_{j},$
where the last inequality results from the fact that study $j$ controls FDR at level $\alpha_{j}$. ∎
To provide some intuitions behind the construction of evidence index in Equation (1), we discuss Step 1 under the light of an important special case of testing a single hypothesis. Suppose $m=d=1$ and we are testing $H_{01}$. Given a valid level $\alpha_{1}$ testing procedure $\delta$, the corresponding evidence index:
$e_{11}=(1/\alpha_{1})\cdot\mathbb{I}(\delta\ \text{rejects}\ H_{01})$ (2)
is the most powerful $e-$value that is equivalent to this testing procedure at threshold $1/\alpha_{1}$. The evidence index in Equation (2) is also known as the all-or-nothing bet (Shafer, 2021) against $H_{01}$ and provides another intuitive interpretation of the evidence indices in Equation (1) as follows: if the $j$-th study controls FDR at level $\alpha_{j}$ then $\mathbf{e}_{j}$ are scaled all-or-nothing bets against the null hypotheses in $\mathcal{M}_{j}$ where the scaling $\frac{m_{j}}{\|\bm{\delta}_{j}\|_{0}\vee 1}$ ensures that if all the null hypotheses corresponding to the non-zero bets $\{i\in\mathcal{M}_{j}:e_{ij}\neq 0\}$ are true then the probability that the sum of those non-zero bets exceed $m_{j}/\alpha_{j}$ is at most $\alpha_{j}$. In Section 4.1 we provide another motivation for the evidence indices and show that $\mathbf{e}_{j}$ naturally arise as building blocks of several popular aggregation and derandomization procedures, such as the derandomized knockoff of Ren and Barber (2024) and the $e-$value aggregation procedure of Li and Zhang (2023).
2.2 Step 2: Evidence Aggregation
IRT aggregates the evidence indices $\mathbf{e}_{j}$ across the studies as follows:
${e}_{i}^{\texttt{agg}}=\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{j=1}^{d}e_{ij}\cdot\mathbb{I}(i\in\mathcal{M}_{j}),\quad\forall i\in\mathcal{M}.$ (3)
In Equation (3), $e_{i}^{\tt agg}$ represents the aggregated evidence across all studies that test hypothesis $H_{0i}$. When each study tests all the $m$ hypotheses and $m_{j}=m$, then $e_{i}^{\tt agg}$ is the arithmetic mean of the $d$ evidence indices corresponding to hypothesis $i$, which essentially dominates any symmetric aggregation function by Proposition 3.1 of Vovk and Wang (2021). However, when $m_{j}$ are different, the aggregation scheme in Equation (3) is a natural counterpart to the arithmetic mean of the $d$ evidence indices. Furthermore, in this setting Theorem 2 establishes that $\mathbf{e}^{\texttt{agg}}=\{e_{i}^{\texttt{agg}}\}_{i\in\mathcal{M}}$ is a set of generalized $e-$values associated with $\mathcal{H}_{0}$.
Theorem 2.
Suppose that each study $j$ controls FDR at level $\alpha_{j}$. Then, $\mathbf{e}^{\texttt{agg}}$ is a set of generalized $e-$values associated with $\mathcal{H}_{0}$.
Proof of Theorem 2..
We have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{\texttt{agg}}]$ $=\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{i\in\mathcal{H}_{0}}\sum_{j=1}^{d}\dfrac{m_{j}}{\alpha_{j}}\cdot\mathbb{E}\left[\dfrac{\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}\right]\cdot\mathbb{I}(i\in\mathcal{M}_{j})\leq\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{j=1}^{d}\dfrac{m_{j}}{\alpha_{j}}\cdot{\rm FDR}(\bm{\delta}_{j})\leq m,$
which completes the proof. ∎
In Section 3.2 we discuss an alternative evidence aggregation scheme under some additional assumptions on the data generating process for each study.
2.3 Step 3: FDR control
In the context of multiple testing with $e-$values, Wang and Ramdas (2022) have proposed the $e-$BH procedure that controls the FDR at level $\alpha$ even under unknown arbitrary dependence between the generalized $e-$values. The $e-$BH procedure is related to the well-known Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) and can be summarized as follows: let $\mathbf{e}=\{e_{i}\}_{i\in\mathcal{M}}$ be a set of generalized $e-$value associated with the null hypotheses $\mathcal{H}_{0}$. Denote $e_{(1)}\geq\ldots\geq e_{(m)}$ as the ordered $e-$values from largest to smallest. The rejection rule of $e-$BH is given by $\delta_{i}=\mathbb{I}\left(e_{i}\geq{m}/{\alpha k_{\alpha}}\right)$ for all $i\in\mathcal{M}$, where the threshold is chosen as
$k_{\alpha}=\max\left\{i\in\mathcal{M}:e_{(i)}\geq\dfrac{m}{i\alpha}\right\}.$ (4)
In step 3, IRT uses $\mathbf{e}^{\tt agg}$ as an input for the $e-$BH procedure to get the final rejection set and, in conjunction with Theorem 2 and Theorem 2 of Wang and Ramdas (2022), the IRT framework controls FDR at level $\alpha$ under unknown arbitrary dependence between the aggregated evidence indices $\mathbf{e}^{\tt agg}$. We summarize the aforementioned three steps in Algorithm 1.
Algorithm 1 Integrative Ranking and Thresholding (IRT)
Inputs: $\mathcal{D}_{j}=\{\bm{\delta}_{j},\alpha_{j},\mathcal{M}_{j}\}$ for $j=1,\ldots,d$ studies, designated FDR level $\alpha\in(0,1)$.
1:// (Step 1) Evidence construction
2:Construct $e_{ij}={m_{j}\delta_{ij}}/\{\alpha_{j}(\|\bm{\delta}_{j}\|_{0}\vee 1)\}$ for each $(j,i)\in[d]\times\mathcal{M}_{j}$.
3:// (Step 2) Evidence aggregation
4:Calculate $e_{i}^{\tt agg}$ as in Equation (3) or ${e}_{i}^{\tt agg*}$ as in Equation (7) if conditions of Theorem 3 hold for each $i\in\mathcal{M}$.
5:// (Step 3) FDR control with $e-$BH procedure
6:if conditions of Theorem 3 hold then
7:     Choose the threshold as $k_{\alpha}=\max\left\{i\in\mathcal{M}:e^{\tt agg*}_{(i)}\geq\dfrac{m}{i\alpha}\right\}$and reject all $H_{0i}$ with $e^{\tt agg*}_{i}\geq m/(\alpha k_{\alpha})$.
8:else
9:     Choose the threshold as $k_{\alpha}=\max\left\{i\in\mathcal{M}:e^{\tt agg}_{(i)}\geq\dfrac{m}{i\alpha}\right\}$and reject all $H_{0i}$ with $e^{\tt agg}_{i}\geq m/(\alpha k_{\alpha})$.
10:end if
3 IRT: more powerful evidence aggregation via multiplication
While the IRT procedure guarantees valid FDR control, the choice of $\alpha$ bears important consideration as far as its power is concerned. For instance, with a relatively smaller value of $\alpha$, IRT may fail to recover discoveries identified by studies with a smaller weight $w_{j}$. In fact, when inferences are pooled with the goal of achieving higher reliability then often $\alpha<\min_{j\in[d]}\alpha_{j}$, and in such settings IRT may exhibit no power. In Section 3.1 we take a simple example to discuss the impact that $\alpha$ has on the power of IRT. Thereafter, in Section 3.2 we present IRT, which relies on an alternative evidence aggregation scheme and is more powerful than IRT when inferences are synthesized for higher reliability.
3.1 Impact of $\alpha$ on the power of IRT
Suppose there are $d=2$ studies, each testing the same set of $m=5$ null hypotheses at levels $\alpha_{1}$ and $\alpha_{2}$, respectively. Consider a simple setting where the $d$ studies reject only the $i$-th null hypothesis. So $\delta_{ij}=1$ and $\|\bm{\delta}_{j}\|_{0}=1$ for all $j\in[d]$. Suppose IRT is used to pool inferences from these studies. The gray shaded region in the left panel of Figure 2 depicts the overall FDR level $\alpha$ required for the $e-$BH procedure to reject $H_{0i}$ when $\alpha_{1}$ varies over $0.001$ to $0.1$ and $\alpha_{2}$ is fixed at $0.05$. Here the red dotted line represents $\min(\alpha_{1},\alpha_{2})$. Most notably, this plot reveals that one must have $\alpha>\min(\alpha_{1},\alpha_{2})$ to reject $H_{0i}$ unless $\alpha_{1}=\alpha_{2}=0.05$, in which case $\alpha$ must at least be $0.05$ for the $e-$BH procedure to reject $H_{0i}$. The right panel considers the same setting but with $d=3,\alpha_{3}=0.03$ and paints a similar picture.
Figure 2: Pooling inferences using IRT. Left: Here $d=2$ studies are testing the same set of $m=5$ null hypotheses at levels $\alpha_{1}$ and $\alpha_{2}$, respectively. Both studies reject only the $i^{\tt th}$ null hypothesis. The gray shaded region depicts the overall FDR level $\alpha$ required for the $e-$BH procedure to reject $H_{0i}$ when $\alpha_{1}$ varies over $0.001$ to $0.1$ and $\alpha_{2}$ is fixed at $0.05$. Here the red dotted line represents $\min(\alpha_{1},\alpha_{2})$. Right: Same setting with $d=3$ and $\alpha_{3}=0.03$.
The calculations for Figure 2 readily follow from the IRT procedure. For instance, in the case of the left panel, IRT rejects $H_{0i}$ at FDR level $\alpha$ if $\alpha\geq 2\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2})$. Furthermore, noting that the quantity on the right side of the inequality is the harmonic mean of $\{\alpha_{1},\alpha_{2}\}$, we have $2\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2})\geq\min(\alpha_{1},\alpha_{2})$ where equality holds when $\alpha_{1}=\alpha_{2}$. However, the results from the left panel are surprising, since intuitively the decisions of the second study should enhance the evidence against $H_{0i}$. How can it be that the evidence from the first study is sufficient to reject $H_{0i}$, but the combined evidence from the two studies is not sufficient to reject $H_{0i}$ unless $\alpha>\min(\alpha_{1},\alpha_{2})$? The reason behind this seemingly counterintuitive situation is that IRT makes no assumption on the dependence between the two studies. For instance, the two studies could be employing identical data and testing methods to reach the same conclusion. Then the collective evidence should not necessarily be interpreted as stronger than the individual evidence from each study, given that the latter study does not introduce any fresh information. For the evidence to truly accumulate, we often presume a certain level of difference in the data or testing procedures applied in these studies. However, such an assumption may be unwarranted when only binary decisions and specified FDR levels are available. Absent any dependence assumptions, the most logical approach would be to remain conservative in our interpretations. The IRT aggregation step described in Section 2.2 follows this principle and aggregates the evidence indices via their arithmetic mean. In fact, it is demonstrated in Vovk and Wang (2021) that without further assumptions on the dependence between $e-$values, the arithmetic mean essentially dominates all other $e-$value aggregation methods.
3.2 Alternative evidence aggregation scheme
If inferences are pooled with the goal of achieving higher reliability then an important implicit assumption is that, informally, the studies are “different” in some sense. In this section we make this idea precise and present IRT, which relies on an alternative evidence aggregation scheme. We will need the following definitions.
Definition 2 (Partial exchangeability).
Let $\{X_{i}\}_{i\in\mathcal{M}}$ be a set of random variables and $\mathcal{I}_{0}$ a subset of $\mathcal{M}$. We say $\mathbf{X}=\{X_{i}\}_{i\in\mathcal{M}}$ is partially exchangeable on $\mathcal{I}_{0}$ if $f(\mathbf{X})=f(\Psi_{i,i^{\prime}}\{\mathbf{X}\})$ for all $i,i^{\prime}\in\mathcal{I}_{0}$, where $\Psi_{i,j}$ is the permutation that swaps the $i$-th and the $j$-th positions, and $f$ is the joint density function of $\mathbf{X}$.
Definition 3 (Symmetric decision rule, Copas (1974)).
A decision rule $\bm{\delta}$ is symmetric if $\bm{\delta}(\Psi\{{\mathbf{X}}\})=\Psi\{\bm{\delta}(\mathbf{X})\}$ for all permutation operators $\Psi$.
Note that the notion of partial exchangeablilty on the set of nulls is commonly used in the conformal inference literature (Bates et al., 2023; Liang et al., 2024) while symmetric decision rules arise naturally in conventional settings where all hypotheses undergo simultaneous testing without the inclusion of auxiliary side information. Lemma 1 guarantees that if for each study the summary statistics are partially exchangeable and the testing procedure is symmetric then a scaled version of the evidence indices in Equation (1) are actually bonafide $e-$values.
Lemma 1.
Suppose for each study $j$ the following holds: (i) the summary statistics $\{X_{ij}\}_{i\in\mathcal{M}_{j}}$ are partially exchangeable on $\mathcal{H}_{0j}$, (ii) the testing procedure is symmetric, and (iii) controls FDR at level $\alpha_{j}$. Then for all $i\in\mathcal{H}_{0j}$, it holds that $\mathbb{E}[e_{ij}]\leq m_{j}/|\mathcal{H}_{0j}|$ with $e_{ij}$ defined in Equation (1). In particular, if $\pi_{j}\in(0,1)$ is a lower bound for $|\mathcal{H}_{0j}|/m_{j}$, then $\pi_{j}e_{ij}$ is an $e-$value, i.e., $\mathbb{E}[\pi_{j}e_{ij}]\leq 1$.
Proof of Lemma 1..
Denote $\theta_{i,i^{\prime}}=(\theta_{i},\theta_{i^{\prime}})$ and ${\theta}_{-i,i^{\prime}}^{j}=\{\theta_{k}\}_{k\neq i,i^{\prime},k\in\mathcal{M}_{j}}$. For all $i,i^{\prime}\in\mathcal{H}_{0j}$, we have
$\mathbb{E}[e_{ij}|\theta_{i,i^{\prime}}=\mathbf{0},\theta_{-i,i^{\prime}}^{j}=\zeta]$ $={\int\mathbb{E}[e_{ij}|\mathbf{X}_{j}]\cdot\mathbb{P}(\mathbf{X}_{j}|\theta_{i,i^{\prime}}=\mathbf{0},\theta_{-i}^{j}=\zeta)\ \mathrm{d}\mathbf{X}_{j}}$
$=\int\mathbb{E}[e_{ij}\mid\boldsymbol{\delta}_{j}(\mathbf{X}_{j})]\cdot{\mathbb{P}(\mathbf{X}_{j}|\theta_{i,i^{\prime}}=\mathbf{0},\theta_{-i}^{j}=\zeta)\ \mathrm{d}\mathbf{X}_{j}}.$
By symmetry of the decision rule we have
$\mathbb{E}[e_{ij}|\bm{\delta}_{j}(\mathbf{X}_{j})]=\mathbb{E}[e_{i^{\prime}j}|\bm{\delta}_{j}(\Psi_{i,i^{\prime}}\{\mathbf{X}_{j}\})].$ (5)
Furthermore, by partial exchangeability we have
$\mathbb{P}(\mathbf{X}_{j}|\theta_{i,i^{\prime}}=\mathbf{0},\theta_{-i,i^{\prime}}^{j}=\zeta)=\mathbb{P}(\Psi_{i,i^{\prime}}\{\mathbf{X}_{j}\}|\theta_{i,i^{\prime}}=\mathbf{0},\theta_{-i,i^{\prime}}^{j}=\zeta).$ (6)
Thus, from equations (5) and (6) we have $\mathbb{E}[e_{ij}]=\mathbb{E}[e_{i^{\prime}j}]$ for all $i,i^{\prime}\in\mathcal{H}_{0j}$. Since $\sum_{i\in\mathcal{H}_{0j}}\mathbb{E}[e_{ij}]\leq m_{j}$ as shown in Theorem 1, we have $\mathbb{E}[e_{ij}]\leq m_{j}/|\mathcal{H}_{0j}|$ for all $i\in\mathcal{H}_{0j}$. ∎
Bonafide $e-$values are substantially more flexible than generalized $e-$values. For instance, $e-$values facilitate reliable inferences for individual hypotheses while generalized $e-$values are limited to simultaneous inference across a set of null hypotheses. Furthermore, the only aggregation scheme for generalized $e-$values discussed in the literature is (weighted) arithmetic mean (Ren and Barber, 2024; Li and Zhang, 2023). In contrast, under independence bonafide $e-$values admit aggregation via multiplication, which allows evidence to “accumulate”.
Let $\mathcal{N}_{i}=\{j:\mathbb{I}(i\in\mathcal{M}_{j})=1\}$ denote the set of studies that test hypothesis $H_{0i}$ with $|\mathcal{N}_{i}|=n_{i}$. Given a pre-determined $k\in\{1,\ldots,n_{i}\}$, denote $\mathcal{S}_{ki}$ as any $k$ element subset of $\mathcal{N}_{i}$, $i\in\mathcal{M}$. Let $\pi\in(0,1)$ be a lower bound of $\min_{j\in[d]}\{|\mathcal{H}_{0j}|/m_{j}\}$ and define $A_{i,\mathcal{S}_{ki}}=\prod_{j\in\mathcal{S}_{ki}}e_{ij}$. The next theorem shows that under certain conditions $e_{i}=\pi^{k}A_{i,\mathcal{S}_{ki}}$ is an $e-$value.
Theorem 3.
Suppose (i) the conditions in Lemma 1 hold and (ii) the summary statistics for the $i$-th testing problem, $\{X_{ij}\}_{j\in\mathcal{N}_{i}}$ are independent conditional on $\theta_{i}$ for all $i$. Let $e_{i}=\pi^{k}A_{i,\mathcal{S}_{ki}}$. Then $\mathbb{E}[e_{i}]\leq 1$ for $i\in\mathcal{H}_{0}$.
Proof of Theorem 3..
Note that $\mathbb{E}\left[A_{i,\mathcal{S}_{ki}}\right]=\prod_{j\in\mathcal{S}_{ki}}\mathbb{E}[e_{ij}]\leq\prod_{j\in\mathcal{S}_{ki}}\dfrac{m_{j}}{|\mathcal{H}_{0j}|}$ and thus $\mathbb{E}[\pi^{k}A_{i,\mathcal{S}_{ki}}]\leq 1$. ∎
Theorem 3 facilitates evidence accumulation via a product rule as we can multiply the $\{e_{ij}\}_{j\in\mathcal{\mathcal{missing}}S_{ki}}$’s for aggregation. However, simply multiplying them may not be ideal since if just one study in $\mathcal{S}_{ki}$ fails to reject $H_{0i}$ the product will be 00. To partially over come this difficulty, we propose to use $e_{i}^{\tt agg*}$ as defined below for evidence aggregation.
$e_{i}^{\tt agg*}=\frac{1}{n_{i}}\sum_{k=1}^{n_{i}}\binom{n_{i}}{k}^{-1}\sum_{\mathcal{S}_{ki}\in\mathcal{B}_{ki}}\pi^{k}A_{i,\mathcal{S}_{ki}},$ (7)
where $\mathcal{B}_{ki}$ is the set of all $k$ element subsets of $\mathcal{N}_{i}$. The idea is to try all possible $\pi^{k}A_{i,\mathcal{S}_{ki}}$ and then take average. Since $\pi^{k}A_{i,\mathcal{S}_{ki}}$ are $e-$values, their average is also an $e-$value. We denote the procedure that applies the $e-$BH procedure on $(e_{1}^{\tt agg*},\ldots,e_{m}^{\tt agg*})$ as IRT. In the real data analysis and numerical experiments of sections 5 and 6 we take $\pi_{j}=0.5$ for all $j\in[d]$, as is often assumed in the literature (Jin and Cai, 2007).
Figure 3: Left: Same setting as the left panel of Figure 2. The gray shaded region depicts the overall FDR level $\alpha$ required for IRT to reject $H_{0i}$ while the green and the gray regions represent what value of $\alpha$ is required by IRT. Right: Here $d=4$ studies are testing the same set of $m=5$ null hypotheses. The first two studies are testing at level $\alpha_{1}$ while the remaining two at $\alpha_{2}$. The first two studies are using exactly the same data and both reject only the $i$-th null hypothesis at FDR level $\alpha_{1}$. Other two studies are independent such that Theorem 3 holds, but both continue to reject the same $i$-th null hypothesis at FDR level $\alpha_{2}$. So $\delta_{ij}=1$ and $\|\bm{\delta}_{j}\|_{0}=1$ for all $j\in[d]$. When IRT is used to pool the inferences, the gray shaded region depicts the overall FDR level $\alpha$ required for the $e-$BH procedure to reject $H_{0i}$ as $\alpha_{1}$ varies and $\alpha_{2}$ is fixed at $0.05$. The gray and the green shaded regions depict what value of $\alpha$ is required for $e-$BH to reject $H_{0i}$ when the hybrid scheme IRT H (see Remark 1) is used.
We now return to the example considered in Section 3.1 but instead use IRT to pool the inferences from the two studies. The green and the gray shaded regions in the left panel of Figure 3 depicts the overall FDR level $\alpha$ required for the $e-$BH procedure to reject $H_{0i}$ as a function of $\alpha_{1}$ and with $\alpha_{2}=0.05$. Clearly, IRT is able to reject $H_{0i}$ even when $\alpha<\min(\alpha_{1},\alpha_{2})$. This represents a stark contrast to IRT which requires $\alpha$ to stay in the gray region to reject $H_{0i}$. The reason for this distinction is related to the fact that, courtesy Theorem 3, the aggregation scheme underlying IRT is more powerful than the arithmetic mean. Note that the calculations for the left panel directly follow from the IRT procedure which, in this example, rejects $H_{0i}$ at FDR level $\alpha$ if $\alpha\geq 8\alpha_{1}\alpha_{2}/(\alpha_{1}+\alpha_{2}+m)$. While this example relies on a relatively simple setting, the numerical experiments in Section 6 confirm the broader conclusion that, in general, IRT is not powerful when $\alpha<\min_{j\in[d]}\alpha_{j}$ while, under the conditions of Theorem 3, IRT is powerful in this setting.
Remark 1.
The preceding discussion naturally begs the question regarding the practical relevance of IRT while pooling inferences for higher reliability. Here, we address this concern.
Motivated by the setting in the left panels of figures 2 and 3, suppose there are $d=4$ studies, each testing the same set of $m=5$ null hypotheses. The first two studies are testing at level $\alpha_{1}$ while the remaining two at $\alpha_{2}$. Consider a setting where the first two studies are using exactly the same data and both reject only the $i$-th null hypothesis at FDR level $\alpha_{1}$. Other two studies are independent such that Theorem 3 holds, but both continue to reject the same $i$-th null hypothesis at FDR level $\alpha_{2}$. So $\delta_{ij}=1$ and $\|\bm{\delta}_{j}\|_{0}=1$ for all $j\in[d]$. To pool inferences across these $d$ studies, IRT and IRT can be used in conjunction. Specifically, the aggregated evidence indices from the first two studies are constructed using IRT, denoted $e_{i}^{\tt agg}$ (Equation (3)) and IRT is employed to aggregate the evidence indices from the remaining two studies, denoted $e_{i}^{\tt agg,*}$ (Equation (7)). Denote $e_{i}^{\tt agg,H}=(1/2)(e_{i}^{\tt agg}+e_{i}^{\tt agg,*})$ as the hybrid aggregated evidence indices. Note that $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are a set of generalized $e-$values under $\mathcal{H}_{0}$. When $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are used as inputs to the $e-$BH procedure, the gray and the green shaded regions in the right panel of Figure 3 depict the overall FDR level $\alpha$ required for $e-$BH to reject $H_{0i}$ as a function of $\alpha_{1}$ and with $\alpha_{2}=0.05$. Clearly, this hybrid scheme is able to reject $H_{0i}$ even when $\alpha<\min(\alpha_{1},\alpha_{2})$. In contrast, the gray shaded region reveals that IRT has no power unless $\alpha>\min(\alpha_{1},\alpha_{2})$.
The above example represents a practical setting where often additional information regarding data-sharing or the use of auxiliary side information for multiple testing is available for the $d$ studies. For instance, suppose prior knowledge dictates that a set of $d_{1}\subset[d]$ studies share data amongst themselves, while the conditions of Theorem 3 hold for the remaining set of $d_{2}=[d]\setminus d_{1}$ studies. Denote the aggregated evidence indices across the $d_{1}$ studies derived from IRT as $\{e_{i}^{\tt agg,d_{1}}\}_{i\in\mathcal{M}}$ and those derived from IRT across the $d_{2}$ studies as $\{e_{i}^{\tt agg*,d_{2}}\}_{i\in\mathcal{M}}$. Then, the hybrid aggregated evidence indices $e_{i}^{\tt agg,H}=(1/d)(|d_{1}|e_{i}^{\tt agg,d_{1}}+|d_{2}|e_{i}^{\tt agg*,d_{2}})$ are a set of generalized $e-$values under $\mathcal{H}_{0}$ and the $e-$BH procedure guarantees FDR control at level $\alpha$ when $\{e_{i}^{\tt agg,H}\}_{i=1}^{m}$ are used as inputs. We call this procedure IRT H and evaluate its empirical performance in Section 6.
4 Discussion
We present additional insights about the IRT framework. In particular, we (i) show that the evidence indices in Equation (1) are connected to some existing aggregation and derandomization procedures (Section 4.1), (ii) prove that IRT guarantees asymptotic FDR control if some studies control their FDR asymptotically (Section 4.2), and (iii) develop a procedure, along the lines of IRT, for pooling inferences when the study-specific $p-$values are available (Section 4.3).
4.1 Connections to existing aggregation and derandomization procedures
Leveraging e-values for aggregation and derandomization for specific FDR methods has been explored in literature recently (Ren and Barber, 2024; Li and Zhang, 2023; Bashari et al., 2023; Zhao and Sun, 2024). In this subsection, we show that these seemly distinct e-values constructions can be viewed as special cases of our construction in Equation (1) when examined from an asymptotic perspective.
A generic FDR procedure can be described abstractly as follows
  1. 1.
    (Ranking) Construct a suitable summary statistics $T_{i}$ for for each $H_{0,i}$ and rank the null hypotheses according to $T_{i}$.
  2. 2.
    (FDP Estimation) For any given $t$ estimate the FDP of the decision rule $\boldsymbol{\delta}(t)=\{\delta_{1}(t),\ldots,\delta_{m}(t)\}$, where $\delta_{i}(t)=\mathds{1}(T_{i}\leq t)$. Denote the estimate as $\widehat{\rm FDP}(t)$.
  3. 3.
    (Thresholding) For a given target FDR level $\alpha$, define $t_{\alpha}=\sup\{t:\widehat{\rm FDP}(t)\leq\alpha\}$. Reject $H_{0,i}$ if and only if $T_{i}\leq t_{\alpha}$.
Denote the above FDR procedure as $\boldsymbol{\delta}=(\delta_{1},\ldots,\delta_{m})$, where $\delta_{i}=\mathbb{I}(T_{i}\leq t_{\alpha})$, then e-values constructed in Ren and Barber (2024); Li and Zhang (2023); Bashari et al. (2023); Zhao and Sun (2024) can be written in the form of
$e_{i}=\dfrac{m\delta_{i}}{\widehat{\rm FDP}(t_{\alpha})\|\boldsymbol{\delta}\|_{0}}.$ (8)
Note that the denominator in Equation (8) can be viewed as the estimated number of false discoveries. In what follows we give two explicit examples showing that the construction in (8) and (1) are equivalent as $\|\boldsymbol{\delta}\|_{0}\rightarrow\infty$.
  1. 1.
    Asymptotic equivalence to Ren and Barber (2024) - The knockoff filter (Barber and Candès, 2015, 2019) is a framework for selecting a set of covariates that are relevant for predicting a response variable $Y$ with guaranteed control of the FDR. For each covariate $X_{i}$, it constructs a statistic $W_{i}$ that is likely to be large if $X_{i}$ is relevant for predicting $Y$ conditional on $\{X_{j}\}_{j\neq i}$ and has a symmetric distribution around 0, otherwise. The knockoff filter selects $X_{i}$ if and only if $W_{i}\geq t_{\alpha}$ where
    $t_{\alpha}=\inf\left\{t>0:\frac{1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t)}{\sum_{i=1}^{m}\mathbb{I}(W_{i}\geq t)}\leq\alpha\right\}.$ (9)
    Ren and Barber (2024) show that the following is a set of generalized $e-$values.
    $e_{i}=\frac{m\cdot\mathbb{I}(W_{i}\geq t_{\alpha})}{1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t_{\alpha})},\quad\forall i\in\mathcal{M},$ (10)
    We now explain how Equation (10) is related to Equation (1). Denote $\delta_{i}=\mathbb{I}(W_{i}\geq t_{\alpha})$. Then $t_{\alpha}$ can be written as
    $t_{\alpha}=\inf\left\{t>0:1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t)\leq\alpha\|\bm{\delta}\|_{0}\right\}.$
    Let $0<\breve{t}_{\alpha}<t_{\alpha}$ be such that $\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-\breve{t}_{\alpha})=1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t_{\alpha})$. We then have
    $\alpha\|\bm{\delta}\|_{0}=\alpha\sum_{i=1}^{m}\mathbb{I}(W_{i}\geq t_{\alpha})\leq\alpha\sum_{i=1}^{m}\mathbb{I}(W_{i}\geq\breve{t}_{\alpha})<1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-\breve{t}_{\alpha})=2+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t_{\alpha}),$
    where the second inequality follows form the definition of $t_{\alpha}$ and $\breve{t}_{\alpha}$. For the decision rule $\delta_{i}=\mathbb{I}(W_{i}\geq t_{\alpha})$ the evidence index defined in Equation (1) becomes ${m\mathbb{I}(W_{i}\geq t_{\alpha})}/{(\alpha\|\bm{\delta}\|_{0})}$ and
    $\frac{m\mathbb{I}(W_{i}\geq t_{\alpha})}{\alpha\|\bm{\delta}\|_{0}}$ $\leq\frac{m\mathbb{I}(W_{i}\geq t_{\alpha})}{1+\sum_{i=1}^{m}\mathbb{I}(W_{i}\leq-t_{\alpha})}\leq\frac{m\mathbb{I}(W_{i}\geq t_{\alpha})}{\alpha\|\bm{\delta}\|_{0}-1}=\frac{m\mathbb{I}(W_{i}\geq t_{\alpha})}{\alpha\|\bm{\delta}\|_{0}}\cdot\frac{\|\bm{\delta}\|_{0}}{\|\bm{\delta}\|_{0}-\alpha^{-1}},$
[…]
  2. 2.
    Asymptotic equivalence to Li and Zhang (2023) - Given null hypotheses $H_{01},\ldots H_{0m}$ and $p-$values $p_{1},\ldots,p_{m}$, the BH procedure with target FDR level $\alpha$ rejects $H_{0i}$ if and only if
    $\delta_{i}=\mathbb{I}(p_{i}\leq t_{\alpha}),\text{~{}where~{}}t_{\alpha}=\sup\left\{t\in(0,1]:\frac{mt}{\sum_{i=1}^{m}\mathbb{I}(p_{i}\leq t)}\leq\alpha\right\},$ (11)
    Li and Zhang (2023) show that the following is a set of generalized $e-$values.
    $e_{i}=t_{\alpha}^{-1}\mathbb{I}(p_{i}\leq t_{\alpha}),\quad\forall i\in\mathcal{M}.$ (12)
    Define
    $\breve{t}_{\alpha}=t_{\alpha}+\frac{1}{m}.$ (13)
    We then have
    $\alpha\|\boldsymbol{\delta}\|_{0}=\alpha\sum_{i=1}^{m}\mathbb{I}(p_{i}\leq t_{\alpha})\leq\alpha\sum_{i=1}^{m}\mathbb{I}(p_{i}\leq\breve{t}_{\alpha})<m\breve{t}_{\alpha}=mt_{\alpha}+1,$ (14)
    where the first and second inequality follows from the definition of $t_{\alpha}$ and $\breve{t}_{\alpha}$ in equations (11) and (13). For the BH procedure, the evidence index in Equation (1) takes the form $m\mathbb{I}(p_{i}\leq t_{\alpha})/(\alpha\|\boldsymbol{\delta}\|_{0})$. Note that
    $\frac{m\mathbb{I}(p_{i}\leq t_{\alpha})}{\alpha\|\bm{\delta}\|_{0}}\leq\frac{m}{mt_{\alpha}}\mathbb{I}(p_{i}\leq t_{\alpha})\leq\frac{m\mathbb{I}(p_{i}\leq t_{\alpha})}{\alpha\|\bm{\delta}\|_{0}-1}=\frac{m\mathbb{I}(p_{i}\leq t_{\alpha})}{\alpha\|\bm{\delta}\|_{0}}\cdot\frac{\|\bm{\delta}\|_{0}}{\|\bm{\delta}\|_{0}-\alpha^{-1}},$
[…]
Remark 2.
The ranking statistic employed in Dai et al. (2023a, b) for derandomization can be expressed as $\delta_{i}/(\widehat{\rm FDP}(t_{\alpha})\|\boldsymbol{\delta}\|_{0}).$ This formulation yields the same ranking as (8). However, Dai et al. (2023a, b) do not use $e$-BH for aggregation.
4.2 Asymptotic FDR control
A key requirement for the validity of the IRT procedure is that the study-specific multiple testing procedure controls FDR at their pre-specified level $\alpha_{j}$. Theorems 2 and 3 implicitly assume that such an FDR control holds for finite samples, i.e. $\mathbb{E}[\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}/\|\bm{\delta}_{j}\|_{0}\vee 1]\leq\alpha_{j}$ for all $j\in[d]$. In reality, however, for some studies their FDR control may be asymptotic in $m_{j}$. In such a scenario, the IRT procedure in Algorithm 1 guarantees FDR control at level $\alpha$ as $m_{j}\to\infty$. We summarize the above discussion in the following proposition.
Proposition 1.
Suppose study $j$ controls FDR at level $\alpha_{j}$ asymptotically, i.e., ${\rm FDR}(\bm{\delta}_{j})\leq\alpha_{j}+o_{p}(1)$. Then, Algorithm 1 ensures that FDR is controlled at level $\alpha$ asymptotically.
Proof of Proposition 1..
We first establish that $\mathbf{e}^{\tt agg}$ in Equation (3) is a set of generalized $e-$values asymptotically. Let $e_{i}^{\tt agg}$ be as defined in Equation (3). Akin to the proof of Theorem 2, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{\texttt{agg}}]$ $=\dfrac{m}{\sum_{j=1}^{d}m_{j}}\sum_{i\in\mathcal{H}_{0}}\sum_{j=1}^{d}\dfrac{m_{j}}{\alpha_{j}}\cdot\mathbb{E}\left[\dfrac{\delta_{ij}}{\|\bm{\delta}_{j}\|_{0}\vee 1}\right]\cdot\mathbb{I}(i\in\mathcal{M}_{j})$
$\leq\dfrac{m}{\sum_{i=1}^{m}m_{j}}\sum_{j=1}^{d}\frac{m_{j}}{\alpha_{j}}\cdot(\alpha_{j}+o_{p}(1))\leq m(1+o_{p}(1)).$
Next, we consider $e_{i}^{\tt agg*}$ from Equation (7). Similarly, following the same arguments in Lemma 1, we can show that $\mathbb{E}[e_{ij}]\leq\pi_{j}(1+o_{p}(1))$. Thus, using the same notation as in Theorem 3, we have $\mathbb{E}[\pi^{k}A_{i,\mathcal{S}_{ki}}]\leq 1+o_{p}(1)$. It follows that $\mathbf{e}^{\tt agg*}=\{e_{i}^{\tt agg*}\}_{i\in\mathcal{M}}$ is also a set of asymptotic generalized $e-$values. Let $\bm{\delta}=\{\delta_{1},\ldots,\delta_{m}\}$ be the decision of $e-$BH applied on a set of asymptotic generalized $e-$values $\mathbf{e}^{\tt agg}$. Note that $\delta_{i}=1$ indicates that $e_{i}^{\texttt{agg}}\geq m/(\alpha\|\bm{\delta}\|_{0}\vee 1)$ based on the decision rule of $e-$BH procedure. Thus, it holds that
$\text{FDR}(\bm{\delta})$ $=\mathbb{E}\left[\sum_{i=1}^{m}\dfrac{\mathbb{I}(\delta_{i}=1,\theta_{i}=0)}{\|\bm{\delta}\|_{0}\vee 1}\right]\leq\mathbb{E}\left[\sum_{i=1}^{m}\dfrac{\alpha}{m}\cdot e_{i}^{\texttt{agg}}\mathbb{I}(\delta_{i}=1,\theta_{i}=0)\right]$
$\leq\frac{\alpha}{m}\cdot\mathbb{E}\left[\sum_{i=1}^{m}e_{i}^{\tt agg}\mathbb{I}(\theta_{i}=0)\right]=\frac{\alpha}{m}\cdot m(1+o_{p}(1))=\alpha+o_{p}(1),$
where the first inequality results from $\mathbb{I}(X>t)\leq X/t$ holds for all non-negative random variable $X$. ∎
4.3 When original $p-$values are available
While the IRT framework takes $\mathcal{D}_{j}$ as an input from each study $j$, it is important to consider the setting where study-specific $p-$values, denoted by $\{p_{ij}\}_{j\in\mathcal{N}_{i},i\in\mathcal{M}}$, are available. The goal under this setting is to aggregate the $p-$values pertaining to each hypothesis $i$ and then determine an appropriate threshold for FDR control at level $\alpha$ using the aggregated $p-$values. However, choosing an aggregation function for combining multiple $p-$values is challenging without making additional assumptions regarding their dependence structure (Vovk and Wang, 2020). Furthermore, if the underlying model is misspecified the validity of the corresponding $p-$values may be affected. In contrast, $e-$values are relatively more robust to such model misspecification (Wang and Ramdas, 2022) and particularly to dependence between the $p-$values (Vovk and Wang, 2021).
In this section, we consider the following calibrator from Equation (B.1) in Vovk and Wang (2021) for transforming $p_{ij}$ into corresponding $e-$values:
$e_{ij}^{\texttt{P2E}}(\kappa)=\begin{cases}\infty~{}&\text{if~{}}p_{ij}=0\\ \dfrac{\kappa(1+\kappa)^{\kappa}}{p_{ij}(-\log p_{ij})^{1+\kappa}}~{}&\text{if~{}}p_{ij}\in(0,e^{-\kappa-1}]\\ 0~{}~{}&\text{if~{}}p_{ij}\in(e^{-\kappa-1},1]\end{cases},$ (15)
where we choose $\kappa=1$ throughout the paper following the recommendation, and write $e_{ij}^{\texttt{P2E}}:=e_{ij}^{\texttt{P2E}}(1)$. Note that $e_{ij}^{\texttt{P2E}}$ as defined in Equation (15) are bonafide $e-$values. Therefore, to aggregate $e_{ij}^{\texttt{P2E}}$ we can simply take the average
${e}_{i}^{\texttt{{P2E},agg}}=\dfrac{1}{n_{i}}\sum_{j=1}^{d}e_{ij}^{\texttt{P2E}}~{}\mathbb{I}(i\in\mathcal{M}_{j}).$ (16)
Furthermore, if the $p-$values $\{p_{ij}\}_{j\in\mathcal{N}_{i}}$ are independent given $\theta_{i}=0$ then, in the spirit of Equation (7), we can aggregate $e_{ij}^{\texttt{P2E}}$ through multiplication as follows:
${e}_{i}^{\texttt{{P2E},agg*}}=\frac{1}{n_{i}}\sum_{k=1}^{n_{i}}\binom{n_{i}}{k}^{-1}\sum_{\mathcal{S}_{ki}\in\mathcal{B}_{ki}}\Big{[}\prod_{j\in\mathcal{S}_{ki}}e_{ij}^{\texttt{P2E}}\Big{]}.$
We denote the method that applies the $e-$BH procedure on ${e}_{i}^{\texttt{{P2E},agg}}$ and $*{e}_{i}^{\texttt{{P2E},agg*}}$ as P2E and P2E, respectively. In Sections 5 and 6 we evaluate the numerical performance of P2E and P2E vis-a-vis IRT and IRT.
Remark 3.
[…]
We note that for study $j$, the $\{e_{ij}\}$’s defined in Equation (1) share the same value whenever $\delta_{ij}=1$. In contrast, $e_{ij}^{\texttt{P2E}}$ retains some information about the magnitude of $p_{ij}$. For instance, if $p_{ij}$ is extremely small, in the sense that $H_{0i}$ will be rejected by study $j$ at a much smaller FDR level than $\alpha_{j}$, it is possible for $e_{ij}^{\texttt{P2E}}$ to be much larger than $e_{ij}$. Table 1 presents a setting where two studies test the same five null hypotheses at target FDR level $0.05$. The hypothetical $p-$values obtained by the two studies are given in the first and fourth rows of Table 1.
Table 1: A summary of the $p-$values and $e-$values by the two hypothetical studies, with the first $p-$value observed by study 1 extremely small.
Hypothesis
1 2 3 4 5
Study 1 $p-$value observed 0.001 0.02 0.05 0.08 0.2
IRT $e-$value 50 50 0 0 0
P2E $e-$value 235.76 6.53 4.46 3.92 0
Study 2 $p-$value observed 0.015 0.02 0.08 0.4 0.6
IRT $e-$value 50 50 0 0 0
P2E $e-$value 7.56 6.53 3.92 0 0
Aggregated IRT evidence index 50 50 0 0 0
Aggregated P2E evidence index 121.66 6.53 4.19 1.96 0
Clearly, the aggregated P2E evidence index (Equation (16)) in Table 1 substantially exceeds the aggregated IRT evidence index (Equation (3)). Furthermore, using the five aggregated P2E evidence indices as inputs, the $e-$BH procedure successfully rejects the first null hypothesis at a significance level of $\alpha=0.045$. In contrast, the $e-$BH procedure with the five aggregated IRT evidence indices fails to make any rejection at level $\alpha$. However, unless the observed $p-$values are exceptionally small relative to the designated FDR threshold, IRT can outperform P2E. Table 2 presents another example. In this setting if the average IRT $e-$value is used as input for the $e-$BH procedure, we are still able to reject the first two null hypotheses at level 0.05. In contrast, if we use the average P2E $e-$value as input for the $e-$BH procedure, none of the five null hypotheses can be rejected at level 0.05.
Table 2: A summary of the $p-$values and $e-$values by the two hypothetical studies.
Hypothesis
1 2 3 4 5
Study 1 $p-$value observed 0.01 0.02 0.05 0.08 0.2
IRT $e-$value 50 50 0 0 0
P2E $e-$value 9.43 6.53 4.46 3.92 0
Study 2 $p-$value observed 0.015 0.02 0.08 0.4 0.6
IRT $e-$value 50 50 0 0 0
P2E $e-$value 7.56 6.53 3.92 0 0
Aggregated IRT evidence index 50 50 0 0 0
Aggregated P2E evidence index 8.50 6.53 4.19 1.96 0
In fact, the numerical experiments in Section 6 reveal that IRT is typically more powerful than P2E when $\alpha\geq\max_{j\in[d]}\alpha_{j}$. When the distribution of the test statistic under the null hypothesis is misspecified, the corresponding $p-$values may not be optimal. In such settings, P2E and P2E may fail to provide valid FDR control. We explore this scenario in Section 6 and find that both IRT and IRT are relatively more robust to such sub-optimal $p-$values.
5 Real data illustration
We illustrate the IRT framework for the integrative analysis of $d=8$ microarray studies (Singh et al., 2002; Welsh et al., 2001; Yu et al., 2004; Lapointe et al., 2004; Varambally et al., 2005; Tomlins et al., 2005; Nanni et al., 2002; Wallace et al., 2008) on the genomic profiling of human prostate cancer. The first three columns of Table 3 summarize the $d$ datasets where a total of
Table 3: Summary of the $d=8$ studies and the evidence against each rejected null hypothesis. Here $e_{j}^{+}=\max\{e_{ij}:i=1,\ldots,m_{j}\}$.
$j$ Study $m_{j}$ Sample size $\alpha_{j}$ $\|\bm{\delta}_{j}\|_{0}$ $e_{j}^{+}$
1 Singh et al. (2002) 8,799 102 0.05 2,094 84.04
2 Welsh et al. (2001) 8,798 34 0.01 921 955.27
3 Yu et al. (2004) 8,799 146 0.05 1,624 108.36
4 Lapointe et al. (2004) 13,579 103 0.05 3,328 81.60
5 Varambally et al. (2005) 19,738 13 0.01 282 6999.29
6 Tomlins et al. (2005) 9,703 57 0.01 1,234 786.30
7 Nanni et al. (2002) 12,688 30 0.01 0 0
8 Wallace et al. (2008) 12,689 89 0.05 4,716 53.81
$m=23,367$ unique genes are analyzed with each gene $i$ being profiled by $n_{i}\in[d]$ studies. The left panel of Figure 4 presents a frequency distribution of the $n_{i}$’s where almost $30\%$ of the $m$ genes are analyzed by just one of the $d$ studies while approximately $18\%$ of the genes are profiled by all $d$ studies.
Our goal in this application is to use the IRT framework to construct a rank ordering of the $m$ gene expression profiles for prostate cancer. Such rank ordering is particularly useful when data privacy concerns prevent the sharing of study-specific summary statistics, such as $p-$values, and information regarding the operational characteristics of the multiple testing methodologies used in each study. For study $j$, our data are an $m_{j}\times s_{j}$ matrix of expression values where $s_{j}$ denotes the sample size in study $j$. Each sample either belongs to the control group or the treatment group and the goal is to test whether gene $i$ is differentially expressed across the two groups. Since IRT operates on the binary decision vector $\bm{\delta}_{j}$, we convert the expression matrices from each study to $\bm{\delta}_{j}$ as follows. For each study $j$, we first use the R-package limma (Ritchie et al., 2015) to get the $m_{j}$ vector of raw $p-$values. Thereafter, the BH procedure is applied to these raw $p-$values at FDR level $\alpha_{j}$ (see column five in Table 3) to derive the final decision sequence $\bm{\delta}_{j}$. We note that typically an important intermediate step before computing the $p-$values in each study is to first validate the quality and compatibility of these studies via objective measures of quality assessment, such as Kang et al. (2012). In this application, however, we do not consider such details.
Figure 4: Left: frequency distribution of $n_{i}$’s. Right: heatmap of the $\log$ evidence indices for $100$ randomly sampled genes across the $d$ studies. White indicates genes not analyzed, while shades of brown represent evidence indices of 00, indicating failure to reject.
The sixth column of Table 3 reports the number of rejections for each of these studies and the last column presents the evidence against each rejected null hypothesis in study $j$. It is interesting to see that study 5 (Varambally et al., 2005) receives the highest evidence for its rejected hypotheses, which is not surprising given the large weight $w_{5}$ that each of its relatively small number of rejections receives. In contrast, study 8 (Wallace et al., 2008) has the smallest non-zero evidence which is driven by the largest number of rejections reported in this study. The right panel of Figure 4 presents a heatmap of the $\log$ evidence indices for $100$ randomly sampled genes across the $d$ studies. Here the white shade represents a gene not analyzed by the study while the shade of brown represents an evidence index of 00 which corresponds to a failure to reject the underlying null hypothesis. The heterogeneity across the $d$ studies is evident through the different magnitudes of the evidence indices constructed for each study.
Table 4: Distribution of rejection overlaps across $7$ studies.
Rejection overlap
$j$ $\|\bm{\delta}_{j}\|_{0}$ 1 2 3 4 5 6 8
1 2,094 - 509 1,531 387 7 130 1,029
2 921 509 - 423 108 1 27 324
3 1,624 1,531 423 - 294 7 105 809
4 3,328 387 108 294 - 17 172 970
5 282 7 1 7 17 - 4 8
6 1,234 130 27 105 172 4 - 365
8 4,716 1,029 324 809 970 8 365 -
Table 4 presents the distribution of rejection overlaps across the $d$ studies, with the exception of study 7. For instance, studies 1 and 3 share $1,531$ rejected hypotheses while studies 2 and 5 share just $1$ rejected hypothesis. Also, study 5, which investigates the largest number of genes, has minimal overlap with the other studies as far as its discoveries are concerned.
5.1 Ranking and thresholding using IRT
In this section we aggregate the evidence indices using Equation (3) and threshold the ordered aggregated evidences using the $e-$BH procedure at $\alpha=0.1$. We recall that this thresholding scheme guarantees valid FDR control under unknown and arbitrary dependence between the aggregated evidences.
Figure 5: Left: histogram of log-transformed non-zero aggregated evidences. Right: scatter of top 25 genes, color and shape-coded by the gene analysis frequency across the $d$ studies. The top and bottom figures employ IRT and P2E respectively
The top-left panel of Figure 5 presents a histogram of the log-transformed non-zero aggregated evidence from IRT, while the top-right panel plots the top 25 genes with respect to their aggregated evidence, colored and shape-coded by the number of times the corresponding gene was analyzed across the $d$ studies. Interestingly, the top second and third genes have $n_{i}=2$ and $3$, respectively, suggesting that apart from the number of times a particular null hypothesis is analyzed across the $d$ studies, the magnitude of the study-specific evidence indices also play a key role in the overall ranking. To put this into perspective, the bottom-right panel of Figure 5 presents the top 25 genes with respect to their aggregated evidence from the P2E framework introduced in Section 4.3. In stark contrast to IRT, here the top 25 genes have $n_{i}\geq 7$. Furthermore, both the left and right panels of Figure 5 suggest that $e_{i}^{\texttt{P2E}}$ can be substantially larger in magnitude than $e_{i}^{\tt agg}$ particularly when one of the studies rejects the null hypothesis with an astronomically small $p-$value.
Table 5: Distribution of rejected hypotheses with respect to $n_{i}$ using IRT and P2E at $\alpha=0.1$.
# Rejections $n_{i}=1$ 2 3 4 5 6 7 8
IRT 2,405 23.91% 2.95% 4.53% 1.25% 5.03% 18.04% 15.13% 29.15%
P2E 5,336 16.38% 8.24% 3.32% 3.88% 2.96% 9.22% 20.48% 35.51%
Next, we study the composition of rejected hypotheses from IRT and P2E at $\alpha=0.1$. Table 5 presents the distribution of rejected hypotheses with respect to $n_{i}$ and reinforces the point that for IRT, the evidence weights $w_{j}$ play a key role in the overall ranking.
5.2 Ranking and thresholding using IRT
Here we aggregate the evidence indices using the scheme discussed in Section 3.2 and threshold the ordered aggregated evidences using the $e-$BH procedure at $\alpha=0.005$. We note that this thresholding scheme guarantees valid FDR control under exchangeability of the study-specific summary statistics (Definition 2), symmetry of the study-specific decision rule (Definition 3) and independence of the $n_{i}$ summary statistics for each testing problem.
Figure 6: Left: histogram of log-transformed non-zero aggregated evidences. Right: scatter of top 25 genes, color and shape-coded by the gene analysis frequency across the $d$ studies. The top and bottom figures employ IRT and P2E respectively.
Table 6: Distribution of rejected hypotheses with respect to $n_{i}$ using IRT and P2E at $\alpha=0.005$.
# Rejections $n_{i}=1$ 2 3 4 5 6 7 8
IRT 472 0 0 2.33% 1.27% 0.63% 48.52% 22.46% 24.79%
P2E 4,129 4.94% 5.38% 1.98% 0.75% 1.40% 12.88% 26.23% 46.43%
The top row of Figure 6 presents a histogram of the log-transformed non-zero aggregated evidence from IRT (top left panel) and a plot of the top 25 genes ranked according to their aggregated evidence, colored and shape-coded by the number of times the corresponding gene was analyzed across the $d$ studies (top right panel). The bottom row presents the same plots for P2E. We find that IRT includes three genes, ranked $1^{\tt st},6^{\tt th},23^{\tt rd}$ with $n_{i}\leq 6$ amongst the top 25. Furthermore, P2E includes two genes, ranked $12^{\tt th},22^{\tt nd}$, with $n_{i}\leq 6$. While this comparison is not as drastic as Figure 5, it continues to suggest that for IRT the magnitude of study-specific evidence indices play an important role in the overall ranking. In contrast, P2E relies on $n_{i}$ and the magnitude of the $p-$values for ranking the $m$ genes. This distinction is further emphasized in Table 6 where IRT rejects an overall higher percentage of hypotheses than P2E when $n_{i}\leq 6$.
6 Numerical experiments
In this section we assess the empirical performances of IRT, IRT and IRT H on simulated data. We consider eight simulation scenarios with $m=1000$ and test $H_{0i}:\mu_{i}=0~{}vs~{}H_{1i}:\mu_{i}\neq 0$, where
$\mu_{i}\stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}0.8\cdot\delta_{\{0\}}+0.1\cdot\mathcal{N}(3,1)+0.1\cdot\mathcal{N}(-3,1),$
and $\delta_{\{a\}}$ denotes a point mass at $a$. In each scenario, study $j$ uses data $X_{ij}$, to be specified subsequently, to conduct $m_{j}$ tests and reports the corresponding decisions $\bm{\delta}_{j}$ obtained from the BH procedure with control level $\alpha_{j}$. The empirical performances of IRT and its derivatives are compared against three alternative procedures: (i) Fisher, which pools the study specific $p-$values using Fisher’s method (Fisher, 1948), and then applies the BH procedure on the pooled $p-$value sequence for FDR control, (ii) the P2E procedure from Section 4.3 and (iii) the P2E procedure, also from Section 4.3. For IRT H, we use the following scheme across all our simulation settings: the IRT aggregation scheme (Equation (3)) is employed for the first $d_{1}=\lfloor d/2\rfloor$ studies and the IRT aggregation scheme (Equation (7)) is used for the remaining $d_{2}=d-d_{1}$ studies. When the null distribution of the test statistic is correctly specified and the corresponding $p-$values are independent, we expect Fisher to exhibit higher power than $e-$value procedures, such as IRT and P2E. Nevertheless, in such settings Fisher provides a practical benchmark for assessing the empirical performances of P2E, P2E, IRT, IRT and IRT H, where the last three procedures rely only the binary decision sequences $\bm{\delta}_{j}$.
Figure 7: FDP and ETP comparisons for Scenario 1.
Scenario 1 (Independent studies). We begin by considering $d$ independent studies. Specifically, we let $X_{ij}\mid\mu_{i},\sigma_{j}\stackrel{{\scriptstyle\rm ind.}}{{\sim}}\mathcal{N}(\mu_{i},\sigma_{j}^{2}),~{}\sigma_{j}\stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}{\rm Unif}(0.75,2),m_{j}=m$, $\alpha_{j}=0.01$, $\alpha=0.005$ and vary $d$ from $5$ to $10$. Figure 7 presents the average FDP and the ETP across 200 Monte Carlo repetitions. We make several observations. First, while all methods control the FDR at $\alpha$, Fisher, unsurprisingly, has the highest power across all values of $d$ and is followed by P2E. Second, IRT H is more powerful than IRT. In fact, the latter exhibits no power since $\alpha<\alpha_{j}$ for all $j\in[d]$ in this setting, further reinforcing the discussion in Section 3.1 and Remark 1. However, this is not the case when the studies are correlated and $\alpha>\max_{j\in[d]}\alpha_{j}$, as scenarios 3 and 4 demonstrate. Third, IRT is less powerful than P2E. The explanation of this phenomenon lies in the observation that in this setting P2E employs optimal $p-$values as the underlying null distribution of the test statistic is fully known. However, when the null distribution is misspecified P2E and P2E may not even provide valid FDR control. Scenarios 7 and 8, reported later in this section, consider such a setting. Finally, IRT is more powerful than IRT H. This is expected since in this setting the conditions of Theorem 3 hold and the aggregation scheme of Equation (7) results in a more powerful procedure than IRT H.
Figure 8: FDP and ETP comparison for Scenario 2.
Scenario 2 (Correlated studies I). The data are generated according to Scenario 1 with $\sigma_{j}=1,d=10,\alpha_{j}=0.01,\alpha=0.005$ but we introduce correlation across the first $d_{1}=\lfloor d/2\rfloor$ studies. In particular, we generate $(X_{i1},\ldots,X_{i\lfloor d/2\rfloor})$ from a multivariate normal distribution and set $\text{Corr}(X_{ij},X_{ik})=\rho$ for all $(j,k)\in\{1,\ldots,\lfloor d/2\rfloor\}$, $j\neq k$, where $\rho\in\{-0.1,0,0.1,0.3,0.5,0.7,0.9\}$. Figure 8 reports the average FDP and the ETP for various methods. In this scenario, the first $d_{1}$ test statistics and the corresponding $p-$values for each hypothesis are not independent unless $\rho=0$, thus violating the conditions of Theorem 3. Consequently, P2E, IRT and Fisher no longer guarantee FDR control. Indeed, the left panel of Figure 8 reveals that P2E and Fisher fail to control the FDR at $0.5\%$ for large $\rho$ and therefore do not appear in the plot for some values of $\rho$. While IRT appears to control the FDR, it has no theoretical support for FDR guarantee in this setting. From the right panel, we find that IRT H is the most powerful procedure in this setting that also provably controls FDR.
Figure 9: FDP and ETP comparison for Scenario 3.
Scenario 3 (Correlated studies II). We continue to borrow the setting from Scenario 1 with $\sigma_{j}=1,~{}d=10,~{}\alpha_{j}=0.01,~{}\alpha=0.05$ and introduce correlation across all $d$ studies. Specifically, we let $\text{Corr}(X_{ij},X_{ik})=\rho$ for all $j\neq k$, where $\rho\in\{-0.1,0,0.1,0.3,0.5,0.7,0.9\}$. Figure 9 reports the average FDP and the ETP for various methods. In this scenario, the $d$ test statistics and the corresponding $p-$values for each hypothesis are not independent unless $\rho=0$. Thus, P2E, IRT, IRT H and Fisher have no theoretical support for FDR control at level $\alpha$. As observed in Scenario 2, P2E and Fisher continue to fail to control the FDR at $5\%$ for large $\rho$ and therefore do not appear in the left panel of Figure 9 for some values of $\rho$. In fact IRT is the most powerful procedure in this setting that also provably controls FDR. While several methods for pooling exchangeable $p-$values are available (see for instance Vovk and Wang (2020)), all of them rely on some prior knowledge regarding the strength of the dependence between the $p-$values, which may not be available in practice.
Scenario 4 (Correlated studies and dependent test statistics). We generate the data from Scenario 1 with $\sigma_{j}=1,\alpha_{j}=0.01,\alpha=0.02$ and introduce correlation across the studies as well as the test statistics. In particular, we let $\text{Corr}(X_{ij},X_{ik})=0.7,~{}j\neq k$ and $\text{Corr}(X_{ij},X_{rj})=0.5,i\neq r$. In this setting too P2E, IRT, IRT H and Fisher no longer enjoy theoretical guarantees for FDR control. Figure 10 reports the average FDP and the ETP for various methods as $d$ varies from $5$ to $10$. We see a similar pattern as Figure 9 where P2E and Fisher fail to control the FDR at $2\%$ while IRT controls the FDR and exhibits higher power than P2E.
Figure 10: FDP and ETP comparison for Scenario 4.
Figure 11: FDP and ETP comparison for Scenario 5.
Scenario 5 (Varying $m_{j}$ and $\alpha_{j}$). In this scenario we go back to the setting of independent studies. The data are generated according to Scenario 1 with $\sigma_{j}=1,m=1000,d=10$ and $\alpha=0.03$, but we vary $(m_{j},\alpha_{j})$ for the $d$ studies. To vary $m_{j}$, we set $m_{(1)}=\max\{m_{1},\ldots,m_{d}\}=900$ and consider the ratio $\eta=\min\{m_{1},\ldots,m_{d}\}/m_{(1)}$. For a given choice of $\eta$, we first sample $m_{1},\ldots,m_{d}$ uniformly from $[\lceil m_{(1)}\eta\rceil,m_{(1)}]$ with replacement and then for each $j$, $m_{j}$ hypotheses are chosen at random from the $m$ hypotheses without replacement. We set $\alpha_{j}\in\{0.05,0.03,0.01\}$ according to $m_{j}\leq 600$, $m_{j}\in(600,800]$ or $m_{j}>800$, respectively. Thus, in this setting studies with a higher $m_{j}$ receive a larger weight on their rejections. Figure 11 reports the average FDP and the ETP for various methods as $\eta$ varies over $[0.1,0.5]$. We find that both IRT H and IRT exhibit higher power as $\eta$ increases and dominate IRT in power for all values of $\eta$. Furthermore, when $\eta$ is large, $m_{j}$ is large and studies receive a relatively higher weight $w_{j}$ on their rejections, which leads to an improved power in this setting.
Figure 12: FDP and ETP comparison for Scenario 6.
Scenario 6 (Sensitivity to $\alpha$). In this scenario we report the performance of the six procedures for varying $\alpha$. The data are generated according to Scenario 1 with $\sigma_{j}=1,m=1000,d=10$, but with $(m_{j},\alpha_{j})$ varying across the $d$ studies. To vary $m_{j}$, we first sample $m_{1},\ldots,m_{d}$ uniformly from $\{2000,2500,3000\}$ with replacement and then set $\alpha_{j}\in\{0.05,0.03,0.01\}$ according to $m_{j}=2000$, $2500$ or $3000$, respectively. Thus, in this setting $m<m_{j}$ and studies with a higher $m_{j}$ continue to receive a larger weight on their rejections. Finally, we uniformly sample the $m$ hypotheses without replacement from $\{1000,\ldots,3000\}$ which ensures that $n_{i}\geq 4$ for all $i\in\{1,\ldots,m\}$. Figure 12 reports the average FDP and the ETP for various methods as $\alpha$ varies over $\{0.001,0.003,0.005,0.01,0.03,0.05,0.1\}$. Several observations are in order. First, the left panel reveals that all methods control FDR at level $\alpha$. Second, the right panel reveals that IRT and IRT H are powerful even when $\alpha<\min_{j\in[d]}\alpha_{j}$ while IRT exhibits power only for $\alpha\geq 0.05$, thus exemplifying the discussion in Section 3.1. Finally, when $\alpha\geq 0.05$ IRT is more powerful than P2E as observed in scenarios 3 and 4.
Figure 13: FDP and ETP comparisons for Scenario 7.
Scenario 7 (misspecified null distribution: I). We consider two settings where the null distribution of the test statistic is misspecified and the corresponding $p-$values are no longer optimal. For setting 1 we let $m_{j}=m,\alpha_{j}=0.05,\alpha=0.01$ and $d=10$. For agent $j$, the summary statistics $X_{ij}=\mu_{i}+Y_{ij}$ where $Y_{ij}$ are i.i.d. central $t-$distributed random variables with $\nu>2$ degrees of freedom. To test $H_{0i}:\mu_{i}=0~{}vs~{}H_{1i}:\mu_{i}\neq 0$, the $p-$values are calculated assuming a standard Normal distribution for the scaled summary statistics, i.e, $p_{ij}=2\Phi(-|X_{ij}|/\sigma)$ where $\sigma=\sqrt{\nu/(\nu-2)}$. Finally, all procedures rely on these $p-$values for inference. Figure 13 reports the average FDP and the ETP for various methods as $\nu$ varies over $\{5,10,15,30,50,100\}$. We find that while Fisher, P2E and P2E fail to control the FDR at $\alpha$ for small $\nu$, IRT and IRT H control the FDR for all values of $\nu$.
Setting 2 borrows the design from Setting 1 but allows the studies to be correlated, i.e, $\text{Corr}(X_{ij},X_{ik})=0.7$ for all $j\neq k$, and sets $\alpha_{j}=0.01,\alpha=0.02$. Since the conditions of Theorem 3 do not hold in this setting and the corresponding $p-$values are not independent, we exclude Fisher, P2E, IRT and IRT H from our comparisons. Figure 13 reports the results of this setting and reveals that P2E does not control the FDR at level $\alpha$ for $\nu=5$. Furthermore, IRT is far more powerful than P2E in this setting for all but the smallest value of $\nu$.
Figure 14: FDP and ETP comparisons for Scenario 8.
Scenario 8 (misspecified null distribution: II). Here we consider two additional settings where the null distribution is misspecified. For setting 1 $m_{j}=m,\alpha_{j}=0.05,\alpha=0.01$ and $d=10$. For agent $j$, the summary statistics $X_{ij}=\mu_{i}+Y_{ij}$ where $Y_{ij}\stackrel{{\scriptstyle i.i.d}}{{\sim}}(1-q)t_{5}+qN(0,1)$ and $q\in[0,1]$. Here $t_{5}$ is a central $t-$distributed random variable with $5$ degrees of freedom. The corresponding $p-$values are calculated assuming a standard Normal distribution for the scaled summary statistics, i.e, $p_{ij}=2\Phi(-|X_{ij}|/\sigma)$ where $\sigma=\sqrt{q+(1-q)5/3}$. Figure 14 reports the average FDP and the ETP for various methods as $q$ varies over $\{0.1,0.3,0.5,0.7,0.9\}$. The three $p-$value based procedures fail to control the FDR at $\alpha$ for all values of $q$. In contrast, IRT and IRT H control the FDR and are also powerful. Setting 2 borrows the design from Setting 1 but allows the studies to be correlated, i.e, $\text{Corr}(X_{ij},X_{ik})=0.7$ for all $j\neq k$, and sets $\alpha_{j}=0.01,\alpha=0.02$. We continue to exclude Fisher, P2E, IRT H and IRT from our comparisons in this setting. Figure 14 reports the results of this setting and reveals that IRT is a relatively better procedure than P2E, both in terms of FDR control and power.
Overall, the results of scenarios 7 and 8 reveal that when the $p-$values are misspecified, procedures that rely directly on the magnitude of the $p-$values for pooling inferences may fail to control FDR at the desired level. In contrast, IRT, IRT H and IRT are relatively robust to such misspecification. In fact in scenarios 7 and 8 they not only control FDR across all levels of misspecification but also retain power. We, however, caution that IRT and its derivatives are not without limitations. A fundamental requirement underlying the proposed procedures is that the studies control FDR at their desired levels. If the study level FDR control is severely violated, either due to misspecified $p-$values or a faulty multiple testing methodology, then our framework may fail to control the overall FDR. Nevertheless, even in those settings the proposed procedures will be relatively less liberal in making false discoveries than methods that rely directly on the magnitudes of misspecified $p-$values for pooling inferences.
7 IRT for alternative forms of Type I error control
Sections 2 and 3 present IRT for integrative FDR control. Here, we extend IRT to alternative forms of Type I error control.
7.1 IRT for $k$-Family-wise error rate ($k$-FWER) control
We consider a setting where study $j$ controls the $k_{j}-$FWER at level $\alpha_{j}$, i.e.
$\mathbb{P}\left(\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}\geq k_{j}\right)\leq\alpha_{j},\quad\forall j\in[d].$ (17)
Under this setting, the three steps of the IRT procedure are as follows.
  1. 1.
    Evidence construction. Suppose $k_{j}>1$. We consider the following evidence index,
    $e_{ij}=\dfrac{m_{j}\delta_{ij}}{\max\{c_{1}\alpha_{j}\|\bm{\delta}_{j}\|_{0},c_{2}(k_{j}-1),\alpha_{j}\}},\ \text{where}~{}c_{1},c_{2}>0~{}\text{and}~{}\frac{1}{c_{1}}+\frac{1}{c_{2}}=1,\quad\forall i\in\mathcal{M}_{j}.$ (18)
    If study $j$ satisfies Equation (17) then the evidence indices in Equation (18) are generalized $e-$values. To see this, let $V_{j}$ denote the number of false rejections made by study $j$. Then,
    $\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}\dfrac{m_{j}\delta_{ij}}{\max\{c_{1}\alpha_{j}\|\bm{\delta}_{j}\|_{0},c_{2}(k_{j}-1),\alpha_{j}\}}\right]$
    $=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}\dfrac{m_{j}\delta_{ij}}{\max\{c_{1}\alpha_{j}\|\bm{\delta}_{j}\|_{0},c_{2}(k_{j}-1),\alpha_{j}\}}\bigg{|}V_{j}>k_{j}\right]\mathbb{P}(V_{j}\geq k_{j})$
    $+\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}\dfrac{m_{j}\delta_{ij}}{\max\{c_{1}\alpha_{j}\|\bm{\delta}_{j}\|_{0},c_{2}(k_{j}-1),\alpha_{j}\}}\bigg{|}V_{j}<k_{j}\right]\mathbb{P}(V_{j}<k_{j})$
    $\leq\dfrac{m_{j}}{c_{1}\alpha_{j}}\alpha_{j}+\frac{m_{j}}{c_{2}}\cdot 1=m_{j}\left(\frac{1}{c_{1}}+\frac{1}{c_{2}}\right)=m_{j}$
    We recommend choosing $c_{1}=(2k_{j}-1)/k_{j},c_{2}=(2k_{j}-1)/(k_{j}-1)$. The rationale is that when $\alpha_{j}\|\bm{\delta}_{j}\|_{0}=k_{j}$ (as both are estimates of the number of false positives), $c_{1}=(2k_{j}-1)/k_{j},c_{2}=(2k_{j}-1)/(k_{j}-1)$ is the solution to the following optimization problem
    $\text{minimize}\ \ \ \\ \max\{c_{1}\alpha_{j}\|\bm{\delta}_{j}\|_{0},c_{2}(k_{j}-1)\}\quad\text{subject to}\quad c_{1},c_{2}>0~{}\text{and}~{}\frac{1}{c_{1}}+\frac{1}{c_{2}}=1.$
    An important special case is $k_{j}=1$. For this setting, we fix $c_{1}=1$ in Equation (18) and recover the evidence index proposed in Equation (1) for integrative FDR control. This is not surprising since any method that controls FWER at level $\alpha_{j}$ also controls FDR at level $\alpha_{j}$.
    We note that if additional information, such as what procedure study $j$ used to control $k_{j}$-FWER, is available then it becomes feasible to devise more powerful generalized $e-$values. For example, if Bonferroni procedure is used (i.e. $H_{0i}$ is rejected by study $j$ if and only if its $p-$value is $\leq k_{j}\alpha_{j}/m_{j}$), then we can verify that
    $e_{ij}=\dfrac{m_{j}\delta_{ij}}{\alpha_{j}k_{j}},\quad\forall i\in\mathcal{M}_{j},$
    is also a generalized $e-$value. To see this, observe that
    $\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}e_{ij}\right]\leq\frac{m_{j}}{k_{j}\alpha_{j}}\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}\right]=\dfrac{m_{j}}{k_{j}\alpha_{j}}\sum_{i\in\mathcal{H}_{0j}}\mathbb{P}(\delta_{ij}=1)\leq\dfrac{m_{j}}{k_{j}\alpha_{j}}\sum_{i\in\mathcal{H}_{0j}}\dfrac{k_{j}\alpha_{j}}{m_{j}}\leq m_{j}.$
  2. 2.
    Evidence aggregation. Since the evidence indices in Equation (18) are generalized $e-$values, Equation (3) provides the evidence aggregation scheme in this setting and Theorem 2 guarantees that these aggregated evidences continue to be generalized $e-$values associated with $\mathcal{H}_{0}$. Furthermore, if the conditions of Theorem 3 hold then Equation (7) represents the aggregated evidence indices.
  3. 3.
    $k$-FWER control. Given the generalized $e-$values $\mathbf{e}^{\tt agg}=\{e_{1}^{\tt agg},\ldots,e_{m}^{\tt agg}\}$ from Step 2 above, IRT rejects $H_{0i}$ if and only if $e_{i}^{\tt agg}\geq m/(\alpha k)$. This procedure controls the $k$-FWER at level $\alpha$ since
    $\mathbb{P}\left(\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}^{\tt agg}\geq\frac{m}{\alpha k}\Big{)}\geq k\right)\leq\dfrac{1}{k}\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}^{\tt agg}\geq\dfrac{m}{\alpha k}\Big{)}\right]\leq\dfrac{1}{k}\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\dfrac{e_{i}^{\tt agg}\alpha k}{m}\right]\leq\alpha.$
7.2 IRT for Per-family error rate (PFER) control
Suppose study $j$’s testing procedure controls PFER at level $k_{j}$, that is $\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0j}}\delta_{ij}\right]\leq k_{j}$ for $j\in[d]$.
  1. 1.
    Evidence construction. We consider the evidence index
    $e_{ij}=\dfrac{m_{j}\delta_{ij}}{k_{j}},\quad\forall i\in\mathcal{M}_{j}.$ (19)
    It is then straightforward to check that $\mathbf{e}_{j}=\{e_{ij}\}_{i\in\mathcal{M}_{j}}$ are a set of generalized $e-$value associated with $\mathcal{H}_{0j}$.
  2. 2.
    Evidence aggregation. The evidence indices in Equation (19) are generalized $e-$values. Therefore, Equation (3) continues to provide the evidence aggregation scheme in this setting and Theorem 2 guarantees that these aggregated evidences are generalized $e-$values associated with $\mathcal{H}_{0}$. Furthermore, if the conditions of Theorem 3 hold then Equation (7) represents the aggregated evidence indices.
  3. 3.
    $k$-PFER control. Given the Genralized $e-$values $\mathbf{e}^{\tt agg}=\{e_{1}^{\tt agg},\ldots,e_{m}^{\tt agg}\}$ from Step 2 above, IRT rejects $H_{0i}$ if and only if $e_{i}^{\tt agg}\geq m/k$. This procedure controls the PFER at level $k$ since
    $\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}\geq m/k\Big{)}\right]\leq\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbb{I}\Big{(}e_{i}k/m\geq 1\Big{)}\right]\leq\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}(e_{i}k/m)\right]\leq\frac{k}{m}\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}e_{i}\right]\leq k.$
Remark 4.
The IRT framework can be used for integrative inference even when studies employ different type I error control metrics. Suppose, for instance, that $d_{1}$ studies control $k-$FWER, $d_{2}$ control PFER and the remaining $d_{3}$ studies control FDR at desired levels, where $d_{i}\subset[d],i=1,2,3$, $\cap_{i=1}^{3}d_{i}=\emptyset$ and $\cup_{i=1}^{3}d_{i}=[d]$. Denote $e_{ij}^{d_{1}},j\in d_{1},$ as the evidence indices for the $d_{1}$ studies from Equation (18). Similarly, $e_{ij}^{d_{2}}$ and $e_{ij}^{d_{3}}$ denote, respectively, the evidence indices for studies $d_{2}$ and $d_{3}$ from equations (19) and (1). Then $e_{i}=m\left\{\sum_{r=1}^{3}(\sum_{j\in d_{r}}e_{ij}^{d_{r}}/\sum_{j\in d_{r}}m_{j})\right\},~{}i\in\mathcal{M}$, are a set of generalized $e-$values under $\mathcal{H}_{0}$, and the $e-$BH procedure can be applied to $\{e_{i}\}_{i\in\mathcal{M}}$ if, for example, FDR control is the goal. Furthermore, both IRT and IRT H are also applicable in this setting if prior knowledge regarding data-sharing or the use of auxiliary side information for multiple testing is available for the $d$ studies.
8 Conclusion
IRT is a general framework for fusion learning in multiple testing that operates on the binary decision sequences available from diverse studies and conducts integrative inference on the common parameter of interest. It guarantees an overall FDR control under arbitrary dependence between the aggregated evidence indices as long as the studies control their FDR at the desired levels. Under additional assumptions on the data generating process for each study (see Theorem 3), IRT and IRT H provide more powerful procedures for pooling inferences than IRT, while controlling the FDR at the same level. Furthermore, the numerical experiments in Section 6 suggest that when the $p-$values are misspecified, procedures that rely directly on the magnitude of these $p-$values for pooling inferences may fail to control FDR at the desired level. In contrast, IRT, IRT H and IRT are relatively robust to such misspecification and also retain power. The proposed framework can be extended to other popular forms of Type I error control, such as FWER control, as demonstrated in Section 7. We conclude this article with a discussion on several avenues for further investigation.
[…]
Accumulation bias in meta-analysis - In the context of meta-analysis for $m=1$ hypothesis, Schure and Grünwald (2019) recently demonstrate that due to the accumulating nature of scientific evidence dependencies arise between study size $d$ or meta-analysis timing and the results of the meta-analysis. This dependence introduces bias in the estimates, termed accumulation bias, and invalidates the sampling distribution of the test statistic, which is often assumed to be standard Normal. Consequently, Type I errors are inflated as the underlying $p-$values are sub-optimal. The authors recommend using a framework involving likelihood ratios for constructing valid tests as no valid $p-$valued based procedure is feasible in this setting. Note that likelihood ratios are $e-$values and so are the aggregated evidence indices underlying IRT. Therefore, an interesting extension of our framework will be to evaluate the performance of IRT for meta-analysis of $m\geq 1$ hypotheses in the presence of accumulation bias.
9 Acknowledgments
We thank Professor Aaditya Ramdas for his valuable insights on an earlier version of the manuscript. T. Banerjee’s research was supported in part by the University of Kansas General Research Fund #2003743. B. Gang’s research was supported by the National Natural Science Foundation of China grant 12201123. All errors are our own.